The Linux HOWTO Index

by Greg Hankins, gregh@sunsite.unc.edu v2.9.0, 16 November 1995 This document contains an index to the Linux HOWTOs and mini-HOWTOs, as well as other information about the HOWTO project.

Contents

1 What Are Linux HOWTOs? 1 2 Where Do I Get Linux HOWTOs? 1 3 HOWTO Index 2 4 HOWTO Project Developement 7

4.1 Wanted HOWTOs and mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 HOWTOs Currently Under Contruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 Writing and Submitting a HOWTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5 Copyright 8

1 What Are Linux HOWTOs?

Linux HOWTOs are documents which describe in detail a certain aspect of configuring or using Linux. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET2 HOWTO and the Printing HOWTO. HOWTOs are comprehensive docs - much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. There are several HOWTO formats available: plain text, PostScript, dvi, and HTML. In addition to the HOWTOs, there are a multitude of mini-HOWTOs. These are short, specific HOWTOs, such as the Colour-ls mini-HOWTO. They are only available in plain text format.

2 Where Do I Get Linux HOWTOs?

HOWTOs can be retrieved via anonymous FTP from the following sites:

• sunsite.unc.edu:/pub/Linux/docs/HOWTO
• tsx-11.mit.edu:/pub/linux/docs/HOWTO
as well as many mirror sites <ftp://sunsite.unc.edu/pub/Linux/MIRRORS.html>.

You can also browse HOWTOs <http://sunsite.unc.edu/mdw/HOWTO/> on the World Wide Web. Many

mirror sites <http://sunsite.unc.edu/mdw/hmirrors.html> also mirror the HOWTO WWW files.

sunsite.unc.edu is heavily used, so PLEASE use a mirror site if possible.

HOWTOs are also posted regularly to the newsgroup comp.os.linux.answers, and other appropriate Linux newsgroups.

3 HOWTO Index

The following Linux HOWTOs are currently available: (The JE HOWTO, Printing HOWTO and SCSI HOWTO are only available in plain text format. This is because they have not been converted to the SGML source format.)

• Linux BootPrompt HOWTO, by Paul Gortmaker <Paul.Gortmaker@anu.edu.au>. List of boot time arguments and overview of booting software. Updated 18 August 1995.

• Linux Bootdisk HOWTO, by Graham Chapman <grahamc@zeta.org.au>. How to create a boot/root maintenance disk for Linux. Updated 25 June 1995.

• Linux Busmouse HOWTO, by Mike Battersby <mib@deakin.edu.au>. Information on bus mouse compatibility with Linux. Updated 2 August 1994.

• Linux CDROM HOWTO, by Jeff Tranter <tranter@software.mitel.com>. Information on CD-ROM drive compatibility for Linux. Updated 18 October 1995.

• Linux Commercial HOWTO, by Harald Milz <Harald.Milz@Linux.org>. Listing of commercial software products for Linux. Updated September 1995.

• Linux Cyrillic HOWTO, by Alexander L. Belikoff <abel@wisdom.weizmann.ac.il>. How to configure Linux for use with the Cyrillic characterset. Updated 17 September 1995.

• Linux DOSEMU HOWTO, by Mike Deisher <deisher@dspsun.eas.asu.edu>. HOWTO about the Linux MS-DOS Emulator, DOSEMU. Updated 11 August 1995.

• Linux Danish HOWTO, by Thomas Petersen <petersen@risoe.dk>. How to configure Linux for use with the Danish characterset. Updated 9 March 1995.

• Linux Distribution HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. A list of Linux distributions. Updated 7 June 1995.

• Linux ELF HOWTO, by Daniel Barlow <daniel.barlow@sjc.ox.ac.uk>. How to install and migrate to the ELF binary file format. Updated 13 September 1995.

• Linux Ethernet HOWTO, by Paul Gortmaker <Paul.Gortmaker@anu.edu.au>. Information on Ethernet hardware compatibility for Linux. Updated 27 May 1995.

• Linux Firewall HOWTO , by David Rudder <drig@execpc.com>. How to set up a firewall using Linux. Updated 17 July 1995.

• Linux Ftape HOWTO, by Kai Harrekilde-Petersen <khp@pip.dknet.dk>. Information on ftape drive compatibility with Linux. Updated 2 July 1995 for ftape-2.03b.

• Linux German HOWTO, by Thomas Ploss <plossts@cip.mathematik.uni-stuttgart.de>. Information on using Linux with German-specific features. Updated 29 May 1995.

• Linux HAM HOWTO, by Terry Dawson <terryd@extro.ucc.su.oz.au>. HOWTO configure amateur radio software for Linux. Updated 9 July 1995.

• Linux HOWTO Index, by Greg Hankins <gregh@sunsite.unc.edu>. Index of HOWTO documents about Linux. Updated 16 November 1995.

• Linux Hardware Compatibility HOWTO, by Tawei Wan <frac@pobox.com>. A list of hardware known to work with Linux. Updated 14 November 1995.

• Linux Hebrew HOWTO, by Yair G. Rajwan <yair@hobbes.jct.ac.il>. How to configure Linux for use with the Hebrew characterset. Updated 12 September 1995.

• Linux INFO-SHEET, by Michael K. Johnson <johnsonm@sunsite.unc.edu>. Generic introduction to the Linux operating system. Updated 31 March 1995.

• Linux Installation HOWTO, by Matt Welsh <mdw@sunsite.unc.edu>. How to obtain and install the Linux software. Updated 11 December 1994.

• Linux JE HOWTO, by Hiroo Yamagata <hiyori13@interramp.com>. Information on JE, a set of Japanese language extensions for Linux.

• Linux Kernel HOWTO, by Brian Ward <ward@blah.tu-graz.ac.at>. Upgrading and compiling the Linux kernel. Updated 10 September 1995.

• Linux Keyboard HOWTO, by Andries Brouwer <aeb@cwi.nl>. Information about the Linux keyboard, console and non-ASCII characters. Updated 8 November 1995.

• Linux META-FAQ, by Michael K. Johnson <johnsonm@sunsite.unc.edu>. A listing of Linux sources of information. Updated 9 August 1995.

• Linux MGR HOWTO, by Vincent Broman <broman@Np.nosc.mil>. Information on the MGR graphics interface for Linux. Updated 16 November 1994.

• Linux Electronic Mail HOWTO, by Vince Skahan <vince@halcyon.com>. Information on Linux-based mail servers and clients. Updated 31 March 1995.

• Linux NET-2 HOWTO, by Terry Dawson <terryd@extro.ucc.su.oz.au>. HOWTO configure TCP/IP networking, SLIP, PLIP, and PPP under Linux. Updated 13 August 1995.

• Linux NIS HOWTO, by Erwin Embsen <erwin@nioz.nl>. Information on using NIS/YP on Linux systems. Updated 24 January 1995.

• Linux News HOWTO, by Vince Skahan <vince@halcyon.com>. Information on USENET news server and client software for Linux. Updated 31 March 1995.

• Linux PCI HOWTO, by Michael Will <Michael.Will@student.uni-tuebingen.de>. Information on PCI-architecture compatibility with Linux. Updated 15 November 1995.

• Linux PCMCIA HOWTO, by Dave Hinds <dhinds@allegro.stanford.edu>. How to install and use PCMCIA Card Services. Updated 17 October 1995.

• Linux PPP HOWTO, by Al Longyear <longyear@netcom.com>. Information on using PPP networking with Linux. Updated 12 July 1995.

• Linux Portuguese HOWTO, by Joao Carlos Rodrigues Pereira <jcrp@caravela.di.fc.ul.pt>. How to configure Linux for use with the Portuguese characterset. Updated 14 September 1995.

• Linux Printing HOWTO, by Grant Taylor <gtaylor@cs.tufts.edu>. HOWTO on printing software for Linux.

• Linux Printing Usage HOWTO, by Matt Foster <mwf@engr.uark.edu>. How to use the printing system for a variety of file types and options. Updated 5 August 1995.

• Linux SCSI HOWTO, by Drew Eckhardt <drew@PoohSticks.ORG>. Information on SCSI driver compatibility with Linux.

• Linux SCSI Programming HOWTO, by Heiko Eissfeldt <heiko@colossus.escape.de>. Information on programming the generic Linux SCSI interface. Updated 14 June 1995.

• Linux Serial HOWTO, by Greg Hankins <greg.hankins@cc.gatech.edu>. Information on use of serial devices and communications software. Updated 9 October 1995.

• Linux Sound HOWTO, by Jeff Tranter <Jeff Tranter@Mitel.COM>. Sound hardware and software for the Linux operating system. Updated 21 October 1995.

• Linux Term HOWTO, by Patrick Reijnen <patrickr@bart.nl>. HOWTO use the ‘term’ communications package on Linux systems. Updated 15 July 1995.

• Linux Tips HOWTO, by Vince Reed <reedv@rpi.edu>. HOWTO on miscellaneous tips and tricks for Linux. Updated 30 August 1994.

• Linux UMSDOS HOWTO, by Jacques Gelinas <jacques@solucorp.qc.ca>. How to install and use the UMSDOS filesystem. Updated 13 november 1995.

• Linux UPS HOWTO, by Harvey J. Stein <hjstein@math.huji.ac.il>. Information on using a UPS power supply with Linux. Updated 16 April 1995.

• Linux UUCP HOWTO, by Vince Skahan <vince@halcyon.com>. Information on UUCP software for Linux. Updated 31 March 1995.

• Linux XFree86 HOWTO, by Matt Welsh <mdw@sunsite.unc.edu>. How to obtain, install, and configure XFree86 3.1.1 (X11R6). Updated 15 March 1995.

The following mini-HOWTOs are available (plain text only):

• Linux Anon FTP Setup mini-HOWTO, by Christopher Klaus <cklaus@shadow.net>. How to set up a secure anonymous FTP site. Updated 24 July 1994.

• Linux Backup with MSDOS mini-HOWTO, by Christopher Neufeld <neufeld@physics.utoronto.ca>. How to backup MSDOS machines with Linux. Updated ???.

• Linux Boca mini-HOWTO, by David H Dennis <david@amazing.cinenet.net>. How to install a Boca 16-port serial card (Boca 2016). Updated ???.

• Linux BogoMips mini-HOWTO, by Wim C.A. van Dorst <baron@clifton.hobby.nl>. Information about BogoMips. Updated 21 September 1995.

• Linux CD Writer mini-HOWTO, by Matt Cutts <cutts@ms.uky.edu>. How to write CDs. Updated 17 December 1994.

• Linux Caching named mini-HOWTO, by Nicolai Langfeldt <janl@ifi.uio.no>. How to set up a caching nameserver. Updated 20 March 1995.

• Linux Colour ls mini-HOWTO, by Thorbjoern Ravn Andersen <ravn@imada.ou.dk>. How to set up the colours with ’ls’. Updated 12 December 1994.

• Linux Consoles mini-HOWTO, by R. Mark Salathiel <RSalathi@nyx.cs.du.edu>. How to set up multiple virtual consoles. Updated 25 December 1994.

• Linux Dip+SLiRP+CSLIP mini-HOWTO, by Zenon Fortuna <zenon@netcom.com:gt;. How to use Dip+SLiRP+CSLIP together. Updated 15 October 1995.

• Linux GUI Development mini-HOWTO, by Philip Markwalder <pgmarkwa@stud.ee.ethz.ch>. Information about GUI Development Software and Books. Updated 6 June 1995.

• Linux getty-ps mini-HOWTO, by Mark Horton <mah@ka4ybr.com>. How to set up ’getty ps’. Updated ???.

• Linux HTML Validation mini-HOWTO, by Keith M. Corbett <kmc@specialform.com>. How to validate HTML 2.0 documents with nsgmls. Updated 29 October 1995.

• Linux IO Port Programming mini-HOWTO, by Riku Saikkonen <rjs@spider.compart.fi>. How to use I/O ports in C programs. Updated 16 November 1995.

• Linux Iomega ZIP mini-HOWTO, by Byron A. Jeff <byron@cc.gatech.edu>. How to use the Iomega ZIP drive with Linux. Updated 24 August 1994.

• Linux Key Setup mini-HOWTO, by Stephen Lee <sl14@cornell.edu>. How to set up cursor control keys. Updated 13 May 1995.

• Linux Keystroke mini-HOWTO, by Zenon Fortuna <zenon@netcom.com>. How to assign special action to keys. Updated 4 April 1995.

• Linux LBX mini-HOWTO, by Paul D. Smith <psmith@BayNetworks.com> How to use Low-Bandwidth X (LBX). Updated 23 October 1995.

• Linux LILO mini-HOWTO, by Cameron Spitzer <cls@truffula.sj.ca.us>. Examples of typical LILO installations. Updated 4 June 1995.

• Linux Large IDE mini-HOWTO, by Patrick LoPresti <patl@lcs.mit.edu>. How to use large IDE drives (kernel versions < 1.1.39). Updated ???.

• Linux Man Page mini-HOWTO, by Jens Schweikhardt <jens@kssun3.rus.uni-stuttgart.de>. How to write man pages. Updated September 1995.

• Linux Modeline mini-HOWTO, by Rick Niles <niles@axp745.gsfc.nasa.gov>. How to use ’modeline’. Updated 11 February 1995.

• Linux Multiple Ethernet mini-HOWTO, by Don Becker <becker@cesdis.gsfc.nasa.gov>. How to use multiple ethernet cards. Updated ???.

• Linux Online Support mini-HOWTO, by lilo <TaRDiS@mail.utexas.edu>. Information about the Linux Internet Support Cooperative. Updated 2 January 1995.

• Linux Proxy ARP mini-HOWTO, by Al Longyear <longyear@netcom.com>. A small treatise on the use of Proxy ARP. Updated 5 December 1994.

• Linux Queue Remote Mail mini-HOWTO, by Leif Erlingsson <Leif.Erlingsson@mailbox.swipnet.se>. How to queue remote mail and deliver local mail. Updated 19 September 1995.

• Linux Reading List, by James H. Haynes <haynes@cats.ucsc.edu>. Interesting books pertaining to Linux subjects. Updated 16 February 1995.

• Linux SLIP+proxyARP mini-HOWTO, by Dave Kennedy <davek@melita.com> Using SLIP and Proxy ARP to connect LANs to the net. Updated 4 February 1994.

• Linux Stacker mini-HOWTO, by Corey Sweeney <corey@bbs.xnet.com>. How to mount MSDOS Stacked filesystems. Updated ???.

• Linux Swap Space mini-HOWTO, by H. Peter Anvin <hpa@yggdrasil.com>. How to share swap space between Linux and Windows. Updated 19 June 1995.

• Linux TIA mini-HOWTO, by Irish <irish@eskimo.com>. How to use ’TIA’ with Linux. Updated 6 August 1995.

• Linux Term Firewall mini-HOWTO, by Barak Pearlmutter <bap@scr.siemens.com>. How to use ’term’ over a firewall. Updated 11 September 1995.

• Linux Tiny News mini-HOWTO, by Kent Lewis <klewis@inst.com>. How to use ’INN’ to make a small news spool. Updated 25 May 1995.

• Linux Token Ring mini-HOWTO, by Mike Eckhoff <meckhoff@zaphod.wayne.esu1.k12.ne.us>. How to use token ring cards. Updated 5 July 1995.

• Linux Visual Bell mini-HOWTO, by Alessandro Rubini <rubini@ipvvis.unipv.it>. How to disable audible bells, and enable visual bells. Updated May 1995.

• Linux WordPerfect mini-HOWTO, by Wade Hampton <tasi029@tmn.com>. How to set up SCO WordPerfect for Linux. Updated 17 May 1995.

• Linux X Notebook mini-HOWTO, by Darin Ernst <darin@castle.net>. How to set up X on notebooks with WD90C24 video chips. Updated 10 August 1995.

• Linux Xterminal mini-HOWTO, by Scot W. Stevenson <scot@catzen.gun.de>. How to connect Xterminals to Linux boxen. Updated July 1995.

4 HOWTO Project Developement

4.1 Wanted HOWTOs and mini-HOWTOs

If you are interested in writing about any of these subject, please get in touch with me at

gregh@sunsite.unc.edu.

• GCC/LIBRARY HOWTO. Discussing installation/upgrading of GCC and shared libraries under Linux. How to build shared libraries. Most of the material would come from Mitchum DeSouza’s GCC FAQ.

• Notebook HOWTO or Notebook mini-HOWTO. A guide to notebooks - information about power management, LCD displays, video controlers, pointers to related material (PCMCIA-HOWTO), and most anything pertaining to notebooks.

• If you have an idea for another HOWTO or mini-HOWTO, please get in touch.

4.2 HOWTOs Currently Under Contruction

• The Printing HOWTO is being completely rewritten, and will be superceded by the Printing Usage and Printing Setup HOWTOs.

• The Tips HOWTO is being rewritten.

• A Partitioning mini-HOWTO.

• A Security HOWTO.

• A DNS HOWTO.

• An Info Services HOWTO.

• A SMP HOWTO (multiprocessor support for Linux).

• An IPX HOWTO.

4.3 Writing and Submitting a HOWTO

If you would like to write a Linux HOWTO document, there are a few guidelines that you should follow.

• Get in touch with me first, and let me know you want to work on something. You can reach me

at gregh@sunsite.unc.edu.

• Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• If you are writing a HOWTO, you must use the Linuxdoc-SGML package, available from sunsite.unc.edu:/pub/Linux/utils/text/, to format the HOWTO. This package allows us to produce LaTeX (for dvi and PostScript), plain text, and HTML from a single source document, and was designed by Matt Welsh specifically for the HOWTOs. This also gives all of the HOWTOs a uniform look.

• If you are writing a mini-HOWTO, you can use any form you wish. mini-HOWTOs are meant to be short, free-form HOWTOs on very specific subjects. In general, these should be quick to write and maintain, so it’s not worth bothering with the Linuxdoc-SGML package. This does not mean that the mini-HOWTOs are any less important! Only plain text versions of mini-HOWTOs are distributed.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname), and the current version number and release date of the software.

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an “FAQ” or “Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs as a model. The SGML source to the HOWTOs is available on Linux FTP sites.

• Lastly, be prepared to receive questions and comments about your writing. There are several hundreds of accesses to the HOWTO collection every day from around the world!

After you have written the HOWTO, mail it to me. If you used Linuxdoc-SGML, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTOs on sunsite.unc.edu and posting them to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

Then, all you have to do is send me periodic updates whenever appropriate.

5 Copyright

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu via email, or at +1 404 853 9989.